If she drops it from rest and the dime falls 14.2m, how long in time was the dime in the air
1
Expert's answer
2020-09-10T14:40:29-0400
"t=\sqrt{\frac{2h}{g}}\\t=\sqrt{\frac{2(14.2)}{9.8}}=1.70\ s"